OpenStudy (anonymous):

The lake contains 10^5m^3 if water with 5% pollution by volume. Each day 10^3m^3 of clean water flows into the lake and 10^3^3 of polluted lake water flows out. I cant seem to form the discrete time equation for for level of water pollution

OpenStudy (anonymous):

Pollution in water over time t = 5(10^3)-10^3t Water over time t = 10^5-5(10^3) +10^3t

OpenStudy (anonymous):

@rouault thats none of the available answers tho